# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ IV.1 Microscopic models for fermionic QCLs
It would be interesting to explore microscopic models that realize gapped fermionic QCLs. A natural strategy is to start with existing models for bosonic QCLs and replace the bosons with fermions. Our results show that, if the model continues to fractionalize, there will be a drastic shift of the ground state degeneracy on a torus. Concretely, one possible route to take is to study fermionic realizations of the bosonic quantum dimer Moessner and Raman (2008) and Balents-Fisher-Girvin Balents et al. (2002) models. Both models can be mapped to hardcore boson models on the kagome lattice with a fixed cluster charge per kagome hexagon Isakov et al. (2011). They are at fillings $\nu=1/2$ and $\nu=3/2$ of the kagome unit cell, respectively. Thus the minimal bosonic TO they can realize is $\mathbb{Z}_{2}$ gauge theory, and both models realize this. Suppose that the hardcore boson is replaced with a spinless fermion. Our proof above demonstrates that the minimal TO that can be realized in such a system is $\mathbb{Z}_{4}$ gauge theory. More generally, it will thus be interesting to study fermionic models of charge frustration and search for gapped QCL insulating phases.
Of particular interest in the context of moiré TMD materials is to explore the physics of microscopic models of these systems in an intermediate coupling regime where neither the Coulomb interaction nor the bandwidth is overwhelmingly strong. This can possibly be pursued with Density Matrix Renormalization Group methods Kiely and Chowdhury (2023); Zhou et al. (2024). Some caution is required in reaching conclusions about experiments as there is still some uncertainty in the details of the microscopic models of these systems due to poorly understood effects of lattice relaxation.
@@ IV.2 Phenomenology of QCLs
One could also ask, for instance in the TMD setting, about the physical properties of QCLs, should they occur, and their experimental manifestation. Choosing a filling with many closely degenerate charge ordered states would promote charge fluctuations, possibly allowing us to realize the intervening QCL phase in this experimental setting. The charge order has been imaged directly in Scanning Tunneling Microscopy experiments deep in the insulator Li et al. (2021). If the QCL is present, then as the displacement field used to tune the phase diagram is increased, the charge order will vanish before the transition to the metal. Alternately, since, at rational $\nu$, the charge order will be described by a discrete order parameter, there will be a finite temperature phase transition which could be detected experimentally. If the corresponding transition temperature vanishes within the regime where the ground state is incompressibile, then an intervening QCL phase must exist, as depicted in Fig. 5.
Figure 5: A temperature, $T$, vs. bandwidth, $t/U$, schematic phase diagram showing the possibility of an intervening QCL phase. For small bandwidths a Wigner-Mott insulator is stabilized at a fixed fractional $\nu$; here a cartoon of a possible $\sqrt{3}\times\sqrt{3}$ charge order is displayed. As the bandwidth is increased the transition temperature to this phase, $T_{\mathrm{co}}$ vanishes. However, below a crossover temperature, $T^{*}$, the system continues to be nearly incompressible. Then, following our discussion, a gapped QCL must be stabilized at zero temperature, even though TO cannot be stabilized at finite temperature in $(2+1)$d. As the bandwidth is increased further the gapped QCL will transition into a compressible metal, shown by $T^{*}$ going to zero. The metal’s Fermi surface is shown below the axis.
@@
We could also ask about the effects of doping the QCL. A simple possibility is that the cheapest charged excitation in the insulator is the $e$-particle which will then have a non-zero density at a small doping. If the $e$-particle is a boson, an allowed possibility in the minimal TO, then a natural doped state is a condensate of $e$, which is a superconductor that breaks the global $\mathrm{U}(1)$ symmetry. The $e$ condensation will completely destroy all the anyons in the theory so that, as is known from other contexts, this state is smoothly connected to a BCS superconductor though it is stabilized by a different mechanism. Note that this superconductor will be fully gapped as it evolves out of a gapped insulator and will have odd parity pairing (as this is the only possibility for spinless fermions).
If, instead the $e$-particle is a fermion, a natural doped state is a non-Fermi liquid metal where the charge carriers are ‘orthogonal’ to the physical electrons, known as an ‘orthogonal metal’ Nandkishore et al. (2012). If the electron filling is changed to $\nu=\frac{p}{q}+\delta$, the doped $e$-particles are at a filling $q\delta$ due to their fractional charge. Thus the $e$-particles will form a Fermi surface with an area that is $q$ times larger than what would be inferred from the density of doped electrons. The increased Fermi momentum may be measured in a local experiment by looking for $2K_{F}$ signatures (Friedel oscillations) in the charge density around impurities. The fractional charge of the carriers in this orthogonal metal can also be directly evidenced through shot noise measurements.
Finally, it is possible that despite the fractionalization in the QCL, the cheapest charged excitation is actually just the electron. Then upon doping the charge will just go in through doping electrons on top of the TO of the QCL. The resulting state will have a conventional Fermi surface of electrons with an area set by the doping density in the usual way. However the TO of the QCL will remain unchanged, and we will have a fractionalized Fermi liquid Senthil et al. (2003) (FL) that violates the conventional Luttinger theorem.
While other possibilities e.g., a first order transition to a conventional Fermi liquid, are not ruled out these observations show that a QCL, if stabilized at the rational filling, can also lead to fascinating physics in a nearby doping range.
@@ IV.3 Theoretical questions on minimal TO
On the more theoretical front, one might also be able to apply the proof strategy we used to determine the minimal TO of other symmetry groups. We highlight a few situations where this question is pertinent. Electronic quantum Hall bilayers at a total Landau level filling $\nu=1/2+1/2$ have long been studied, both in experiment and in theory. The standard fate is the development of interlayer coherent order (an exciton condensate) accompanied by an integer quantum Hall effect. Refs. Sodemann et al. (2017); Potter et al. (2017) discussed an alternate symmetric gapped phase, which could be stabilized for some range of microscopic parameters, that has $\mathbb{Z}_{4}$ TO. The global symmetry of this system has two $\mathrm{U}(1)$s for each layer, as well as an anti-unitary symmetry that flips particles to holes in both the layers. Ref. Sodemann et al. (2017) further conjectured that the $\mathbb{Z}_{4}$ TO is minimal (in terms of GSD). A proof of this conjecture is accessible with our techniques Musser et al. .
@@
The quantum Hall bilayer system has an anomaly involving charge conservation and particle-hole symmetries Sodemann et al. (2017); Potter et al. (2017). More generally, the (2+1)$d$ surface state of a Symmetry Protected Topological (SPT) phase in (3+1)$d$ has an anomalous symmetry realization and cannot be trivially gapped. In a large class of states, a symmetric gapped state with TO is however possible. What then is the minimal TO for a system with a global symmetry $G$ that has an anomaly?
@@
Another quantum Hall bilayer system where the minimality question arose recently is discussed in the paper Jian et al. (2024) which considered a system with a quantized fractional spin Hall response. The symmetry group is then a product of two $\mathrm{U}(1)$s for each spin/valley along with time reversal that interchanges them Wen and Potter (2024). With time reversal, the authors of Jian et al. (2024) were able to show that the minimal order according to quantum dimension was also $\mathbb{Z}_{4}$ gauge theory at a fractional spin Hall coefficient of one half [^13] Though the $\mathbb{Z}_{4}$ gauge theory appears in both problems, there is a slight difference in the action of the anti-unitary symmetry in this system and the Quantum Hall bilayer system considered in Refs. Sodemann et al. (2017); Potter et al. (2017). However, their proof leaves open the possibility that there might exist a non-Abelian SET with a smaller anyon count but larger quantum dimension. We suspect that some variant of the proofs used here can rule out such a possibility. Indeed, the proofs here should be useful to demonstrate the minimal size of the subset of Abelian anyons in any SET which respects an order two anti-unitary symmetry.
@@
Finally, the concept and definition of minimal order raises some other interesting theoretical questions. First, to what extent should we expect minimal order to be realized in a physical system respecting a symmetry group $G$? Ultimately whatever order is stabilized will depend sensitively on energetics. A recent example is illustrated by the numerical identification of non-minimal non-Abelian quantum spin Hall states in twisted MoTe2 Abouelkomsan and Fu (2024); Jian et al. (2024). In this situation minimal order only seems to be energetically favored if pairing between spin species is sufficiently strong Zhang (2024a, b). However, it might nonetheless be the case that minimal order is generically favored in other situations, e.g. if the interactions are short-ranged in real space. Secondly, one could ask if there are other sensible classes of minimal order and what their relationship to $mG_{\mathcal{D}}$ and $mG_{\rm GSD}$ is. Finally, one could ask if it is possible to pick any TO in SETG and “flow” to a minimal order by successive condensation of anyons which do not break the symmetries $G$. If this is true then minimal orders could be regarded as the “fixed points” of all TOs in SETG.
@@ Acknowledgements.
We thank Chao-Ming Jian for a discussion about Jian et al. (2024), Fiona Burnell for a discussion about the possibility of non-Abelian background anyons, and Arkya Chatterjee, Mike Crommie, Hart Goldman, Ho Tat Lam, Kin Fai Mak, Salvatore Pace, Zhengyan “Darius” Shi, Daniel Spiegel, Ashvin Vishwanath, and Liujun Zou for their willingness to discuss the work and suggest references. We also thank Zhihuan Dong, Nandini Trivedi, Ruben Verresen, and Mike Zaletel for helpful discussions of a related upcoming work. S.M. is supported by the Laboratory for Physical Sciences. M.C. was supported by NSF grant DMR-1846109. T.S. was supported by NSF grant DMR-2206305, and partially through a Simons Investigator Award from the Simons Foundation. This work was also partly supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651446, T.S.).
@@ Appendix A Results about minimal classes
Let SETG be the set of all SETs that respect a symmetry $G$. Then define:
$mG_{\mathcal{D}}=\{\mathcal{C}\in\text{SET}_{G}\mid\mathcal{D}_{\mathcal{C}}\leq\mathcal{D}_{\mathcal{C}^{\prime}}\ \forall\mathcal{C}^{\prime}\in\text{SET}_{G}\},$ (11)
@@
$mG_{\rm GSD}=\{\mathcal{C}\in\text{SET}_{G}\mid|\mathcal{C}|\leq|\mathcal{C}^{\prime}|\ \forall\mathcal{C}^{\prime}\in\text{SET}_{G}\},$ (12)
@@
Suppose that there exists some SET $\mathcal{C}\in mG_{\rm GSD}$ such that $\mathcal{A}=\mathcal{C}$, i.e. $\mathcal{C}$ is Abelian. Then every SET in $mG_{\mathcal{D}}$ is Abelian. In fact $mG_{\mathcal{D}}$ is precisely the subset of Abelian TOs in $mG_{\rm GSD}$.
@@
Consider $\mathcal{C}^{\prime}\in mG_{\mathcal{D}}$, by definition we know that $\mathcal{D}_{\mathcal{C}^{\prime}}^{2}\leq\mathcal{D}_{\mathcal{C}}^{2}=|\mathcal{C}|$ since $\mathcal{C}$ is Abelian. Since $|\mathcal{C}^{\prime}|\leq\mathcal{D}_{\mathcal{C}^{\prime}}^{2}$ we see that $|\mathcal{C}^{\prime}|\leq\mathcal{D}_{\mathcal{C}^{\prime}}^{2}\leq|\mathcal{C}|$. But since $\mathcal{C}\in mG_{\rm GSD}$ we must have $|\mathcal{C}|\leq|\mathcal{C}^{\prime}|$. These inequalities are thus equalities. This means $\mathcal{D}_{\mathcal{C}^{\prime}}^{2}=|\mathcal{C}^{\prime}|$, so $\mathcal{C}^{\prime}$ is an Abelian TO. It also means that $|\mathcal{C}^{\prime}|=|\mathcal{C}|$, so $\mathcal{C}^{\prime}\in mG_{\rm GSD}$ as well.
On the other hand if $\mathcal{C}^{\prime}\in mG_{\rm GSD}$ is Abelian, then we see that $\mathcal{D}_{\mathcal{C}^{\prime}}^{2}=|\mathcal{C}^{\prime}|\leq|\mathcal{C}^{\prime\prime}|\leq\mathcal{D}_{\mathcal{C}^{\prime\prime}}^{2}$ for every $\mathcal{C}^{\prime\prime}\in SET_{G}$. Thus $\mathcal{C}^{\prime}\in mG_{\mathcal{D}}$. We therefore conclude that $mG_{\mathcal{D}}$ is precisely the subset of Abelian TOs in $mG_{\rm GSD}$. ∎
@@
Note that an obvious corrollary of this is that if $mG_{\rm GSD}$ contains only Abelian TOs, then $mG_{\rm GSD}=mG_{\mathcal{D}}$.
@@
The opposite direction is not true, i.e. it may be the case that $mG_{\mathcal{D}}$ contains an Abelian SET, but every SET in $mG_{\rm GSD}$ is non-Abelian. An example of this is fermionic TOs with $\mathrm{U}(1)_{f}$ symmetry, with the requirement that the Hall condutivity $\sigma_{H}=1/2$. The set $mG_{\mathcal{D}}$ contains multiple theories including the Moore-Read state, $\mathrm{U}(1)_{8}$ and others, all with $\mathcal{D}^{2}=16$. However, one can show that all theories in $mG_{\rm GSD}$ have anyon count 12, saturated by the Moore-Read state. Moreover, as the next proof will make clear, all of these theories must be non-Abelian.
@@
Suppose that there exists some SET $\mathcal{C}\in mG_{\mathcal{D}}$ such that $\mathcal{D}_{\mathcal{C}}^{2}>|\mathcal{C}|$, i.e. $\mathcal{C}$ is non-Abelian. Then every SET in $mG_{\rm GSD}$ is non-Abelian.

### v2

@@ IV.1 Microscopic models for fermionic QCLs
It would be interesting to explore microscopic models that realize gapped fermionic QCLs. A natural strategy is to start with existing models for bosonic QCLs and replace the bosons with fermions. Our results show that, if the model continues to fractionalize, there will be a drastic shift of the ground-state degeneracy on a torus. Concretely, one possible route to take is to study fermionic realizations of the bosonic quantum dimer Moessner and Raman (2008) and Balents-Fisher-Girvin Balents et al. (2002) models. Both models can be mapped to hardcore boson models on the kagome lattice with a fixed cluster charge per kagome hexagon Isakov et al. (2011). They are at fillings $\nu=1/2$ and $\nu=3/2$ of the kagome unit cell, respectively. Thus the minimal bosonic TO they can realize is $\mathbb{Z}_{2}$ gauge theory, and both models realize this. Suppose that the hardcore boson is replaced with a spinless fermion. Our proof above demonstrates that the minimal TO that can be realized in such a system is $\mathbb{Z}_{4}$ gauge theory. More generally, it will thus be interesting to study fermionic models of charge frustration and search for gapped QCL insulating phases.
Of particular interest in the context of moiré TMD materials is to explore the physics of microscopic models of these systems in an intermediate coupling regime where neither the Coulomb interaction nor the bandwidth is overwhelmingly strong. This can possibly be pursued with density-matrix renormalization-group methods Kiely and Chowdhury (2023); Zhou et al. (2024). Some caution is required in reaching conclusions about experiments because there is still some uncertainty in the details of the microscopic models of these systems due to poorly understood effects of lattice relaxation.
@@ IV.2 Phenomenology of QCLs
One could also ask, for instance in the TMD setting, about the physical properties of QCLs, should they occur, and their experimental manifestation. Choosing a filling with many closely degenerate charge ordered states would promote charge fluctuations, possibly allowing us to realize the intervening QCL phase in this experimental setting. The charge order has been imaged directly in scanning tunneling microscopy experiments deep in the insulator Li et al. (2021). If the QCL is present, then as the displacement field used to tune the phase diagram is increased, the charge order will vanish before the transition to the metal. Alternately, since, at rational $\nu$, the charge order will be described by a discrete order parameter, there will be a finite-temperature phase transition which could be detected experimentally. If the corresponding transition temperature vanishes within the regime where the ground state is incompressible, then an intervening QCL phase must exist, as depicted in Fig. 5.
Figure 5: A temperature $T$ vs bandwidth $t/U$ schematic phase diagram showing the possibility of an intervening QCL phase. For small bandwidths a Wigner-Mott insulator is stabilized at a fixed fractional $\nu$; here a cartoon of a possible $\sqrt{3}\times\sqrt{3}$ charge order is displayed. As the bandwidth is increased, the transition temperature to this phase, $T_{\mathrm{co}}$, vanishes. However, below a crossover temperature, $T^{*}$, the system continues to be nearly incompressible. Then, following our discussion, a gapped QCL must be stabilized at zero temperature, even though TO cannot be stabilized at finite temperature in $(2+1)$d. As the bandwidth is increased further the gapped QCL will transition into a compressible metal, shown by $T^{*}$ going to zero. The metal’s Fermi surface is shown below the axis.
@@
We could also ask about the effects of doping the QCL. A simple possibility is that the cheapest charged excitation in the insulator is the $e$-particle which will then have a nonzero density at a small doping. If the $e$-particle is a boson, an allowed possibility in the minimal TO, then a natural doped state is a condensate of $e$, which is a superconductor that breaks the global $\mathrm{U}(1)$ symmetry. The $e$ condensation will completely destroy all the anyons in the theory so that, as is known from other contexts, this state is smoothly connected to a BCS superconductor though it is stabilized by a different mechanism. Note that this superconductor will be fully gapped because it evolves out of a gapped insulator and will have odd parity pairing (because this is the only possibility for spinless fermions).
If, instead the $e$-particle is a fermion, a natural doped state is a non-Fermi-liquid metal where the charge carriers are “orthogonal” to the physical electrons, known as an “orthogonal metal” Nandkishore et al. (2012). If the electron filling is changed to $\nu=\frac{p}{q}+\delta$, the doped $e$-particles are at a filling $q\delta$ due to their fractional charge. Thus the $e$-particles will form a Fermi surface with an area that is $q$ times larger than what would be inferred from the density of doped electrons. The increased Fermi momentum may be measured in a local experiment by looking for $2K_{F}$ signatures (Friedel oscillations) in the charge density around impurities. The fractional charge of the carriers in this orthogonal metal can also be directly evinced through shot noise measurements.
Finally, it is possible that despite the fractionalization in the QCL, the cheapest charged excitation is actually just the electron. Then upon doping the charge will just go in through doping electrons on top of the TO of the QCL. The resulting state will have a conventional Fermi surface of electrons with an area set by the doping density in the usual way. However, the TO of the QCL will remain unchanged, and we will have a fractionalized Fermi liquid Senthil et al. (2003) (FL) that violates the conventional Luttinger theorem.
While other possibilities, e.g., a first-order transition to a conventional Fermi liquid, are not ruled out, these observations show that a QCL, if stabilized at the rational filling, can also lead to fascinating physics in a nearby doping range.
We note that throughout our discussion we have assumed the electrons are spinless, but if the electron spin is not polarized then there are additional phenomenological possibilities. These include the possibility of phases which are both QCLs and QSLs.
@@ IV.3 Theoretical questions on minimal TO
On the more theoretical front, one might also be able to apply the proof strategy we used to determine the minimal TO of other symmetry groups. We highlight a few situations where this question is pertinent. Electronic quantum Hall bilayers at a total Landau level filling $\nu=1/2+1/2$ have long been studied, both in experiment and in theory. The standard fate is the development of interlayer coherent order (an exciton condensate) accompanied by an integer quantum Hall effect. References Sodemann et al. (2017); Potter et al. (2017) discussed an alternate symmetric gapped phase, which could be stabilized for some range of microscopic parameters, that has $\mathbb{Z}_{4}$ TO. The global symmetry of this system has two $\mathrm{U}(1)$s for each layer, as well as an anti-unitary symmetry that flips particles to holes in both the layers. Reference Sodemann et al. (2017) further conjectured that the $\mathbb{Z}_{4}$ TO is minimal (in terms of GSD). A proof of this conjecture is accessible with our techniques Cheng et al. (2025).
@@
The quantum Hall bilayer system has an anomaly involving charge conservation and particle-hole symmetries Sodemann et al. (2017); Potter et al. (2017). More generally, the $(2+1)d$ surface state of a symmetry protected topological (SPT) phase in $(3+1)d$ has an anomalous symmetry realization and cannot be trivially gapped. In a large class of states, a symmetric gapped state with TO is, however, possible. What then is the minimal TO for a system with a global symmetry $G$ that has an anomaly?
@@
Another quantum Hall bilayer system where the minimality question arose recently is discussed in the paper Jian et al. (2024) which considered a system with a quantized fractional spin Hall response. The symmetry group is then a product of two $\mathrm{U}(1)$s for each spin (equivalently valley) along with time reversal that interchanges them Wen and Potter (2024). With time reversal, the authors of Ref. Jian et al. (2024) were able to show that the minimal order according to quantum dimension was also $\mathbb{Z}_{4}$ gauge theory at a fractional spin Hall coefficient of one half. [^18] Although the $\mathbb{Z}_{4}$ gauge theory appears in both problems, there is a slight difference in the action of the anti-unitary symmetry in this system and the quantum Hall bilayer system considered in Refs. Sodemann et al. (2017); Potter et al. (2017). However, their proof leaves open the possibility that there might exist a non-Abelian SET with a smaller anyon count but larger quantum dimension. We suspect that some variant of the proofs used here can rule out such a possibility. Indeed, the proofs here should be useful to demonstrate the minimal size of the subset of Abelian anyons in any SET which respects an order two anti-unitary symmetry.
@@
Finally, the concept and definition of minimal order raises some other interesting theoretical questions. First, to what extent should we expect minimal order to be realized in a physical system respecting a symmetry group $G$? Ultimately, whatever order is stabilized will depend sensitively on energetics. A recent example is illustrated by the numerical identification of nonminimal non-Abelian quantum spin Hall states in twisted MoTe2 Abouelkomsan and Fu (2024); Jian et al. (2024). In this situation minimal order only seems to be energetically favored if pairing between spin species is sufficiently strong Zhang (2024a, b). However, it might nonetheless be the case that minimal order is generically favored in other situations, e.g., if the interactions are short-ranged in real space. Second, one could ask if there are other sensible classes of minimal order and what their relationship to $mG^{(\nu)}_{\mathcal{D}}$ and $mG^{(\nu)}_{\rm GSD}$ is. Finally, one could ask if it is possible to pick any TO in SETG,ν and “flow” to a minimal order by successive condensation of anyons which do not break the symmetries $G$. If this is true then minimal orders could be regarded as the “fixed points” of all TOs in SETG,ν.
@@ Acknowledgements.
We thank Chao-Ming Jian for a discussion about Ref. Jian et al. (2024), Fiona Burnell for a discussion about the possibility of non-Abelian background anyons, and Arkya Chatterjee, Mike Crommie, Hart Goldman, Ho Tat Lam, Kin Fai Mak, Salvatore Pace, Zhengyan “Darius” Shi, Daniel Spiegel, Ashvin Vishwanath, and Liujun Zou for their willingness to discuss the work and suggest references. We also thank Zhihuan Dong, Nandini Trivedi, Ruben Verresen, and Mike Zaletel for helpful discussions of a related upcoming work. S.M. is supported by the Laboratory for Physical Sciences. M.C. was supported by NSF grant DMR-1846109. T.S. was supported by NSF grant DMR-2206305, and partially through a Simons Investigator Award from the Simons Foundation. This work was also partly supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651446, T.S.).
@@ Appendix A Results about minimal classes
Let SETG,ν be the set of all SETs that respect a symmetry $G$ and fractional topological response $\nu$. Then define:
$mG^{(\nu)}_{\mathcal{D}}=\{\mathcal{C}\in\text{SET}_{G,\nu}\mid\mathcal{D}_{\mathcal{C}}\leq\mathcal{D}_{\mathcal{C}^{\prime}}\ \forall\mathcal{C}^{\prime}\in\text{SET}_{G,\nu}\},$ (11)
@@
$mG^{(\nu)}_{\rm GSD}=\{\mathcal{C}\in\text{SET}_{G,\nu}\mid|\mathcal{C}|\leq|\mathcal{C}^{\prime}|\ \forall\mathcal{C}^{\prime}\in\text{SET}_{G,\nu}\},$ (12)
@@
Suppose that there exists some SET $\mathcal{C}\in mG^{(\nu)}_{\rm GSD}$ such that $\mathcal{A}=\mathcal{C}$ i.e., $\mathcal{C}$ is Abelian. Then every SET in $mG^{(\nu)}_{\mathcal{D}}$ is Abelian. In fact $mG^{(\nu)}_{\mathcal{D}}$ is precisely the subset of Abelian TOs in $mG^{(\nu)}_{\rm GSD}$.
@@
Consider $\mathcal{C}^{\prime}\in mG^{(\nu)}_{\mathcal{D}}$, by definition we know that $\mathcal{D}_{\mathcal{C}^{\prime}}^{2}\leq\mathcal{D}_{\mathcal{C}}^{2}=|\mathcal{C}|$ since $\mathcal{C}$ is Abelian. Since $|\mathcal{C}^{\prime}|\leq\mathcal{D}_{\mathcal{C}^{\prime}}^{2}$ we see that $|\mathcal{C}^{\prime}|\leq\mathcal{D}_{\mathcal{C}^{\prime}}^{2}\leq|\mathcal{C}|$. But since $\mathcal{C}\in mG^{(\nu)}_{\rm GSD}$ we must have $|\mathcal{C}|\leq|\mathcal{C}^{\prime}|$. These inequalities are thus equalities. This means $\mathcal{D}_{\mathcal{C}^{\prime}}^{2}=|\mathcal{C}^{\prime}|$, so $\mathcal{C}^{\prime}$ is an Abelian TO. It also means that $|\mathcal{C}^{\prime}|=|\mathcal{C}|$, so $\mathcal{C}^{\prime}\in mG^{(\nu)}_{\rm GSD}$ as well.
On the other hand if $\mathcal{C}^{\prime}\in mG^{(\nu)}_{\rm GSD}$ is Abelian, then we see that $\mathcal{D}_{\mathcal{C}^{\prime}}^{2}=|\mathcal{C}^{\prime}|\leq|\mathcal{C}^{\prime\prime}|\leq\mathcal{D}_{\mathcal{C}^{\prime\prime}}^{2}$ for every $\mathcal{C}^{\prime\prime}\in SET_{G,\nu}$. Thus $\mathcal{C}^{\prime}\in mG^{(\nu)}_{\mathcal{D}}$. We therefore conclude that $mG^{(\nu)}_{\mathcal{D}}$ is precisely the subset of Abelian TOs in $mG^{(\nu)}_{\rm GSD}$. ∎
@@
Note that an obvious corrollary of this is that if $mG^{(\nu)}_{\rm GSD}$ contains only Abelian TOs, then $mG^{(\nu)}_{\rm GSD}=mG^{(\nu)}_{\mathcal{D}}$.
@@
The opposite direction is not true i.e., it may be the case that $mG^{(\nu)}_{\mathcal{D}}$ contains an Abelian SET, but every SET in $mG^{(\nu)}_{\rm GSD}$ is non-Abelian. An example of this is fermionic TOs with $\mathrm{U}(1)_{f}$ symmetry, with the requirement that the Hall condutivity $\sigma_{H}=1/2$. The set $mG^{(\sigma_{H})}_{\mathcal{D}}$ contains multiple theories including the Moore-Read state, $\mathrm{U}(1)_{8}$ and others, all with $\mathcal{D}^{2}=16$. However, one can show that all theories in $mG^{(\sigma_{H})}_{\rm GSD}$ have anyon count 12, saturated by the Moore-Read state. Moreover, as the next proof will make clear, all of these theories must be non-Abelian.
@@
Suppose that there exists some SET $\mathcal{C}\in mG^{(\nu)}_{\mathcal{D}}$ such that $\mathcal{D}_{\mathcal{C}}^{2}>|\mathcal{C}|$ i.e., $\mathcal{C}$ is non-Abelian. Then every SET in $mG^{(\nu)}_{\rm GSD}$ is non-Abelian.